24.1. Find the Laplace transform Y(s) of the solution to each of the following initial-value problems. Just find Y (s) usingh. y + 5y + 6y = 4t with y(0) = 1 and y(0) = 0)I need help with d and h. Thank you.
Answer #1

@ y - 4y = t², y(o)= 1, y(o)=3 Take Laplace on both sides 2 ly) - 4 L (y) = L(+3) - [5²Y(S)-syco)-yco)]-4 y(s) = 6 s²y(s)-  1 + 5y + 6y = eht, yco) = 1 and y/o)=0 Lly) + 5L (4) + 6L (y) = 2 lett) 5-y(s)- syco)-y (0)] + 5 [sycs)-yco)] + 6ycs) = I. (

*****Your honest feedback is very important for better results*****

Thanks